# Stable long-term evolution in numerical relativity

### v1

@@ I Introduction
A remarkable example of such weak instabilities is the superradiant extraction of mass from a spinning black hole by a massive probe scalar field [8, 9]. In this process, the mass of the scalar field serves as a reflective outer boundary, enabling continuous mass extraction from the black hole, leading to the formation of a scalar cloud encircling it [10]. However, the growth rate of such superradiant instabilities is extremely slow, posing significant numerical challenges for studying the long-term evolution of this phenomenon [11, 12]. To evade this issue, research on the non-linear evolution of these set-ups has shifted to scenarios involving a charged scalar field around a charged black hole or a massive vector field around a spinning black hole, where the instability rates are significantly faster [13, 14]. Despite this advantage, the studies still require robust numerical techniques capable of simulating long-term evolution. It has been shown, for instance, that a saturated cloud of massive vectors can radiate slowly through gravitational wave emission over time-scales extending to about $10^{5}$ in units of the initial black hole mass, as the hole spins down and exits the superradiant regime [14]. In addition to superradiance, weak non-linear instabilities have also been reported in spacetimes featuring a stable light ring [15, 16]. Numerical studies suggest that exotic compact objects may experience light-ring instabilities after long-term evolution [16]. Conversely, simulations indicate that black holes with multiple photon spheres appear free of such instabilities, even after extended evolution [7], although again diagnosing this outcome is only possible by means of long-term simulations.
@@
On the other hand, the BSSN formulation does not propagate Hamiltonian constraint violations. To address this limitation, the so-called conformal covariant Z4 system (CCZ4), and the related Z4c, have been proposed, in which constraint violations are promoted to dynamical variables [23, 24, 25]. The CCZ4 formulation effectively converts the violations of the Hamiltonian and momentum constraints into wave propagation equations. In contrast, the Z4c formulation focuses solely on propagating the Hamiltonian constraint violation, preserving a structure closer to the BSSN formulation [26, 25]. Similarly to the generalized BSSN formulation, the CCZ4 formulation has also been extended to a fully covariant form by introducing an additional reference metric [27, 28, 29]. While the CCZ4 formulation has demonstrated higher accuracy in simulating binary neutron stars compared to the BSSN formulation, it suffers from non-linear numerical instabilities when applied to black hole spacetimes. As shown in previous studies [30, 27], these instabilities can be alleviated by adequately introducing damping for the constraint violations. However, a key observation is that the damping effect must not be excessively strong to avoid the emergence of further numerical instabilities.
@@ II Numerical schemes for evolving the Einstein equations
which governs the evolution of constraint violations ${}^{(4)}Z_{\mu}$. In weakly curved spacetimes, this approximates a standard damped wave equation. Here, $\kappa_{1}$ and $\kappa_{2}$ serve as damping factors aimed at suppressing the amount of violation of the constraints. In the 3+1 decomposition, ${}^{(4)}Z_{\mu}$ is split into components along the normal vector $n^{\mu}$ and directions tangential to the spatial hypersurface $\Sigma$,
@@
Regarding the fixing of gauge, we employ the $1+\textrm{log}$ slicing condition for the lapse field $\alpha$ and the $\Gamma$-driving condition for the shift vector field $\beta^{i}$ [31, 32, 33, 20],
@@
where $C^{i}$ is an auxiliary field and $\eta$ is another damping parameter, and for simplicity we choose $\eta=1$ throughout this paper.
@@ Acknowledgements.
We are grateful to Yiqian Chen and Shenkai Qiao for useful discussions and valuable comments. SGS, GG and XW are supported by the NSFC (Grant Nos. 12250410250 and 12347133). PW is supported in part by the NSFC (Grant Nos. 12105191, 12275183, 12275184 and 11875196).

### v2

@@ I Introduction
A remarkable example of such weak instabilities is the superradiant extraction of mass from a spinning black hole by a massive probe scalar field [8, 9]. In this process, the mass of the scalar field serves as a reflective outer boundary, enabling continuous mass extraction from the black hole, leading to the formation of a scalar cloud encircling it [10]. However, the growth rate of such superradiant instabilities is extremely slow, posing significant numerical challenges for studying the long-term evolution of this phenomenon [11]. It has been reported that for a rapidly rotating black hole, the dominant scalar mode associated with this instability exhibits a minimum growth time-scale of order $10^{6}$ in units of the black hole mass [12, 13]. To evade this issue, research on the non-linear evolution of these set-ups has shifted to scenarios involving a charged scalar field around a charged black hole or a massive vector field around a spinning black hole, where the instability rates are significantly faster [14, 15]. [^1] Besides these technical considerations, charged scalar and vector fields certainly deserve attention due to their physical properties. For instance, hairy black holes formed through the superradiant growth of a charged scalar field are dual to superconductor systems in the context of the AdS/CFT correspondence [16]. Despite this advantage, the studies still require robust numerical techniques capable of simulating long-term evolution. It has been shown, for instance, that a saturated cloud of massive vectors can radiate slowly through gravitational wave emission over time-scales extending to about $10^{5}M$, as the hole spins down and exits the superradiant regime [15]. In addition to superradiance, weak non-linear instabilities have also been reported in spacetimes featuring a stable light ring [17, 18]. Numerical studies suggest that exotic compact objects may experience light-ring instabilities after long-term evolution, with time-scales of order $10^{4}M$ [18]. Conversely, simulations indicate that black holes with multiple photon spheres appear free of such instabilities, even after extended evolution [7], although again diagnosing this outcome is only possible by means of long-term simulations.
@@
On the other hand, the BSSN formulation does not propagate Hamiltonian constraint violations. To address this limitation, the so-called conformal covariant Z4 system (CCZ4), and the related Z4c, have been proposed, in which constraint violations are promoted to dynamical variables [25, 26, 27]. The CCZ4 formulation effectively converts the violations of the Hamiltonian and momentum constraints into wave propagation equations. In contrast, the Z4c formulation focuses solely on propagating the Hamiltonian constraint violation, preserving a structure closer to the BSSN formulation [28, 27]. Similarly to the generalized BSSN formulation, the CCZ4 formulation has also been extended to a fully covariant form by introducing an additional reference metric [29, 30, 31]. While the CCZ4 formulation has demonstrated higher accuracy in simulating binary neutron stars compared to the BSSN formulation, [^2] See also [32, 33] for detailed comparisons of the performance of BSSN, Z4c and CCZ4. it suffers from non-linear numerical instabilities when applied to black hole spacetimes. As shown in previous studies [34, 29], these instabilities can be alleviated by adequately introducing damping for the constraint violations. However, a key observation is that the damping effect must not be excessively strong to avoid the emergence of further numerical instabilities (see e.g. [34] or Fig. 2 below).
@@ II Numerical schemes for evolving the Einstein equations
which governs the evolution of constraint violations ${}^{(4)}Z_{\mu}$. In weakly curved spacetimes, this approximates a standard damped wave equation. In the 3+1 decomposition, ${}^{(4)}Z_{\mu}$ is split into components along the normal vector $n^{\mu}$ and directions tangential to the spatial hypersurface $\Sigma$,
@@
Regarding the fixing of gauge, we employ the $1+\textrm{log}$ slicing condition for the lapse field $\alpha$ and the $\Gamma$-driver condition for the shift vector field $\beta^{i}$ [35, 36, 37, 22],
@@
where $C^{i}$ is an auxiliary field and $\eta$ is another damping parameter, and for simplicity we choose $\eta=1$ throughout the main text of this paper. Modifications to the form of the $\Gamma$-driver condition are discussed briefly in Appendix A.
@@ Acknowledgements.
We are grateful to Yiqian Chen, Cristian Joana and Shenkai Qiao for useful discussions and comments. SGS, GG and XW are supported by the NSFC (Grant Nos. 12250410250 and 12347133). SGS also acknowledges support from a Provincial Grant (Grant No. 2023QN10X389). PW is supported in part by the NSFC (Grant Nos. 12105191, 12275183, 12275184 and 11875196).
Appendix A Modifications to the $\Gamma$-driver gauge condition
The $\Gamma$-driver gauge condition, Eq. (21), has been shown to be important in achieving stable long-term simulations of black holes [36]. Throughout the main text we have used the standard choice $p=3/4$ and $\eta=1$ in the equation for the auxiliary vector,
$\partial_{\perp}C^{i}=p\partial_{\perp}\tilde{\Lambda}^{i}-\eta C^{i}.$ (38)
Both parameters $p$ and $\eta$ are however adjustable (see e.g. [47] for a comparison of different values in another context). While $\eta$ plays the role of a damping parameter, $p$ may be associated to the speed of propagation of the shift vector, at least when the spacetime is approximately flat [36]. In particular, $p=3/4$ means that the longitudinal component of $\beta^{i}$ propagates at the speed of light, which explains why this is a standard choice and also why larger values of $p$ may be prone to numerical instabilities, as indeed we have checked to be the case in our simulations. Moreover, we find that the long-term instability of the BSSN formulation, at least in the set-up of a Schwarzschild black hole, is also worsened by decreasing the parameter $p$ (cf. the top-left panel of Fig. 7). Thus the choice $p=3/4$ appears to be roughly optimal within the context of long-term black hole evolution.
On the other hand, we have found that reducing the damping coefficient $\eta$ may result in improvements in the stability of the BSSN formulation when applied to a Schwarzschild black hole (top-right panel of Fig. 7). Still, BSSN is seen to underperform when compared with the CCZ3 scheme (bottom-left panel), although the long-term instability appears to be resolved for the choice $\eta=0$. However, this seems to be peculiar to the Schwarzschild case, and indeed we can see that already in the set-up of a RN black hole the BSSN formulation with zero damping leads to instabilities in a relatively short timescale (bottom-right panel). Our simulations also suggest that the impact of changing $\eta$ in the CCZ3 scheme is unimportant.
Figure 7: Evolution of Schwarzschild ($Q=0$) and RN ($Q=0.995$) black holes, as measured by the area of the apparent horizon, $A_{h}$, using various parameter settings in the $\Gamma$-driver gauge condition. The top two panels show the persistence of the long-term instability of a Schwarzschild black hole in the BSSN formulation under changes in the coefficients $p$ and $\eta$ in Eq. (38). While the choice $\eta=0$ appears to cure the issue in the Schwarzschild case (bottom-left panel), this choice results in unstable simulations in the charged black hole set-up (bottom-right). In both cases, the CCZ3 scheme is seen to outperform BSSN.